OpenStudy (anonymous):

How would you solve this using symbolic form/reasoning: x+3=10/x

OpenStudy (saifoo.khan):

\[x + 3 = \frac{10}{x}\]Like that right?

OpenStudy (anonymous):

Yeah, like that.

OpenStudy (saifoo.khan):

Multiply "x" on both sides, \[x(x+3) = 10\]\[x^2 + 3x = 10\]

OpenStudy (anonymous):

Ok. Got that.

OpenStudy (saifoo.khan):

Please hit good answer.

OpenStudy (saifoo.khan):

Thanks.

OpenStudy (anonymous):

Your welcome. And Thanks.